OpenStudy (anonymous):

Simplify 2cos^2(x/2)-1

OpenStudy (gorv):

cos(2x)=2cos^2(x)-1 this is actual formula

OpenStudy (gorv):

so replace x by x/2

OpenStudy (gorv):

\[\cos(2*\frac{ x }{ 2 })=2os^2(\frac{ x }{ 2 })-1\]

OpenStudy (gorv):

= cosx

OpenStudy (anonymous):

Hey I know it's been over a week but thanks for the help